OpenStudy (anonymous):

Five weeks after the introduction of a new toy, the manufacturer found that the demand for the toy was modeled by , while the company’s supply capability was modeled by . Here x is the number of weeks the toy has been on the market, and the values of f(x) and g(x) are in millions of toy units. About how many weeks after the toy was introduced did supply equal demand? Round to the nearest whole number of weeks.

OpenStudy (anonymous):

what r the functions?????????

OpenStudy (anonymous):

f(x)=x^2+5x+4/X^2+6x-7 g(x)=-0.89x^2+850x-50/800+70.5

OpenStudy (anonymous):

so u need to find the value of x for that f(x) =g(x) x^2+5x+4/X^2+6x-7=-0.89x^2+850x-50/800+70.5

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well i used a calculator and found x= 2 i dont make the algebraic part that could resuld in a equation of 4 grade after symplify

OpenStudy (anonymous):

wats the answeerrr